OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

are Whole numbers with or without sign?

OpenStudy (anonymous):

with out?

OpenStudy (michele_laino):

for example Whole numbers are in this set? \[\mathbb{N} \] or in this set \[\mathbb{Z}^+ \cup \mathbb{Z}^- \cup {0}\]

OpenStudy (anonymous):

???

OpenStudy (michele_laino):

sorry, what do you means with "Whole numbers"? Usually I mean numbers which belongs to \[\mathbb{N} \]

OpenStudy (anonymous):

What???

OpenStudy (anonymous):

so Whole Numbers can have sign?

OpenStudy (michele_laino):

conjecture: if Whole numbers are numbers which belong to set \[\mathbb{N} \] then your answers are B. and C.